MR Bank has the following balance sheet:

Table 3: MR Bank balance sheet

Assets Liabilities

Required 

Reserves X Deposits Y 

Excess 

Reserves R50 million

The required reserve ratio is 10%. What is the value of MR bank’s deposits?

A. R 50 million.

B. R 55.6 million.

C. R 10 million.

D. R 45 million.


1
Expert's answer
2021-09-03T09:00:53-0400

required reserve of 10%

excess reserve of 50 million

A

excess reserve=actual reserve −required reserve

deposit= 50m+50m=100m

required reserve = required reserve ratio * deposit calculated

= 0.1 * 100 =10 million

B

50m+55.6m=105.6m

required reserve = required reserve ratio * deposit calculated

= 0.1 * 106.6 =10.56 million

C

50m+10m=60m

required reserve = required reserve ratio * deposit calculated

= 0.1 * 60 =6 million

D

45m+50m=95

required reserve = required reserve ratio * deposit calculated

= 0.1 * 95=9.5million